# Private Evolution Converges

### v1

@@ 4.1 Convergence of PE in Euclidean space
We start by presenting the details of the algorithm that we analyze. Consider a metric (data) space $(\Omega,\rho)$ and a private dataset $S\in\Omega^{n}$. Even though we will use the notation $(\Omega,\rho)$ throughout this section, we assume in this subsection that $\Omega\subset\mathbb{R}^{d}$ is compact with $\operatorname{diam}(\Omega)\leq D$ and $\rho(\cdot,\cdot)$ is the $\ell_{2}$ distance. We also assume Algorithm 2 uses the same APIs as Algorithm 1, as described in Section 3.
Algorithm 2 Private Evolution with $D_{BL}$ projection; steps in blue differ from practical PE [31]
@@
Algorithm 2 iteratively refines an initial random dataset $S_{0}$ using DP nearest-neighbor information. At each iteration, it generates variations $V_{t}$ from the current synthetic dataset $S_{t-1}$ and computes a histogram $\hat{\mu}_{t}$ indicating how often each candidate is the nearest neighbor of points in the sensitive dataset $S$. Gaussian noise is added to this histogram to ensure differential privacy, resulting in a noisy vector $\tilde{\mu}_{t}$. Rather than using thresholding like Algorithm 1 in Line 8, this algorithm projects $\tilde{\mu}_{t}$ onto the probability simplex using the $D_{BL}$ distance in Line 8, yielding a valid distribution $\mu_{t}^{\prime}$ over $V_{t}$ that remains close to the noisy estimate under the BL metric. The next synthetic dataset $S_{t}$ is created randomly by sampling $n_{s}$ points from $V_{t}$ according to $\mu_{t}^{\prime}$ in Line 9, as opposed to Algorithm 1 which creates $S_{t}$ deterministically in Line 9. This process is repeated for $T$ iterations, after which the final synthetic dataset $S_{T}$ is returned. Note that unlike Algorithm 1, our version receives the number of synthetic data points as input. We provide a comparison between Algorithms 1 and 2 in Appendix B.
@@
Algorithm 2 is nearly identical to the practical PE implementation in [31], differing only in how the noisy histogram $\tilde{\mu}_{t}$ is post-processed (highlighted in blue). We project $\tilde{\mu}_{t}$—a signed measure from Gaussian noise added to a histogram—onto $\mathcal{P}(V_{t})$ by minimizing the bounded Lipschitz distance, which can be done via linear programming (e.g., Algorithm 2 from [24]). This is crucial for obtaining worst-case convergence guarantees. In contrast, [31] applies a simpler heuristic: thresholding $\tilde{\mu}_{t}$ at some $H>0$ and re-normalizing, which is more efficient but can discard all votes if values fall below the threshold, making it unsuitable for worst-case analysis. In Section 4.2, we provide data-dependent accuracy bounds for a related noisy histogram obtained from adding Laplace noise and thresholding. These bounds are vacuous in the worst case but can be much tighter in favorable regimes, e.g., when the true private data distribution is highly clustered.
@@
We are now ready to present our main result: an upper bound on expected utility of Algorithm 2. We provide a proof sketch, and the full proof is in Appendix D.
@@
Let $(\Omega,\|\cdot\|_{2})$ with $\Omega\subset\mathbb{R}^{d}$ and $\operatorname{diam}(\Omega)\leq D$ be the sample space. Fix $\sigma>0$. Suppose $\alpha=D\sigma^{1/\max\{d,2\}}$ in (1), and let $S_{T}$ be the output of Algorithm 2 run on input $S\in\Omega^{n},T\geq\lceil\log(\gamma/[D\sigma^{1/\max\{d,2\}}])/\gamma\rceil,n_{s}=\sigma^{-1}(2\lceil\log_{2}(D/\alpha)\rceil+1)^{1/\max\{d,2\}-1},\sigma,\rho(\cdot,\cdot)$. Then
@@
Furthermore, if $\varepsilon,\delta\in(0,1)$, and we instead run Algorithm 2 with the same parameter setting as above except $T=\lceil\log(\gamma(n\varepsilon)^{1/\max\{d,2\}}/(4\sqrt{\log(1/\delta)})^{1/\max\{d,2\}})/\gamma\rceil$ and $\sigma=\frac{4\sqrt{T\log(1.25/\delta)}}{n\varepsilon}$, then the algorithm is $(\varepsilon,\delta)$-DP and its output $S_{T}$ satisfies
@@
At a high level, the convergence of PE follows from two facts regarding the terms on the right-hand side of the inequality:
@@
These statements are made rigorous in our proof. Regarding the first statement, Lemmata D.1 and D.2 give, $\mathbb{E}_{t}[W_{1}(\mu_{S},\hat{\mu}_{t+1})]\leq(1-\gamma)W_{1}(\mu_{S},\mu_{S_{t}})+\alpha$ for some $\gamma=\Theta(1/d)$, where $\mathbb{E}_{t}$ denotes the expectation conditioned on the randomness of the algorithm up to iteration $t$.
To bound $D_{BL}(\hat{\mu}_{t+1},\tilde{\mu}_{t+1})$, we use the fact that $\tilde{\mu}_{t+1}=\hat{\mu}_{t+1}+Z$ with $Z\sim\mathcal{N}(0,\sigma^{2}I_{|V_{t}|})$, which by definition implies $D_{BL}(\hat{\mu}_{t+1},\tilde{\mu}_{t+1})=\sup_{f\in{\cal F}_{BL}}\int_{\Omega}f(d\hat{\mu}_{t+1}-d(\hat{\mu}_{t+1}+Z))=\sup_{f\in\mathcal{F}}\sum_{i\in[|V_{t+1}|]}f(V_{t+1}[i])Z_{i}$, which is the supremum of a Gaussian process. The expectation of this term can be bounded using empirical process theory; Lemma D.3 states that $\mathbb{E}_{t}[D_{BL}(\hat{\mu}_{t+1},\tilde{\mu}_{t+1})]\leq|V_{t}|\sigma G_{|V_{t}|}({\cal F}_{BL})$, where $G_{|V_{t}|}({\cal F}_{BL})$ is a Gaussian complexity term [5]. This term can be bounded using Corollary D.1, which is a consequence of Lemma D.4. We use Dudley’s chaining to prove these results. We remark that our technique prove this bound resembles [24], with the difference that they deal with a Laplacian complexity arising from the analysis of a pure-DP algorithm that uses the Laplace mechanism [18].
Finally, to control $W_{1}(\mu^{\prime}_{t+1},\mu_{S_{t+1}})$ we use results from the literature of the convergence of empirical measures in the Wasserstein distance [29, 19] that quantify the Wasserstein distance between a measure and the empirical measure of iid samples from it; see Lemma D.5. We obtain $\mathbb{E}_{t}[W_{1}(\mu^{\prime}_{t+1},\mu_{S_{t+1}})]\leq\tilde{O}\big{(}n_{s}^{-1/\max\{2,d\}}\big{)}$. Putting all the inequalities together, we conclude that
$\mathbb{E}_{t}[W_{1}(\mu_{S},\mu_{S_{t+1}})]\leq(1-\gamma)W_{1}(\mu_{S},\mu_{S_{t}})+\alpha+2|V_{t}|\sigma G_{|V_{t}|}({\cal F}_{BL})+\tilde{O}\big{(}n_{s}^{-1/\max\{2,d\}}\big{)}.$
Recursing this inequality, we can get a bound on $\mathbb{E}[W_{1}(\mu_{S},\mu_{S_{T}})]$ while carefully balancing the number of variations with the number of synthetic samples so that the error terms are as small as possible. This finishes the proof of $\mathbb{E}\left[W_{1}(\mu_{S},\mu_{S_{T}})\right]\leq\tilde{O}(d\sigma^{1/d})$.
Regarding the second result, for privacy we use the fact that $\sigma=\frac{\sqrt{4T\log(1.25/\delta)}}{n\varepsilon}$ is enough to ensure $(\varepsilon,\delta)$-DP of an adaptive composition of $T$ Gaussian mechanisms where each of them adds noise to a NN histogram with $\ell_{2}$ sensitivity $\sqrt{2}/n$ ([14], Corollary 3.3). For convergence, we plug in the expression for $T$ in $\sigma=\frac{\sqrt{4T\log(1.25/\delta)}}{n\varepsilon}$ and then $\sigma$ in $\mathbb{E}\left[W_{1}(\mu_{S},\mu_{S_{T}})\right]\leq\tilde{O}(d\sigma^{1/d})$. ∎
@@
We emphasize that most of our proof techniques work in metric spaces, so under some conditions on the APIs the analysis of Algorithm 2 can be extended to more general Banach spaces than the one considered in our result. See Subsection F for more details.
@@ 4.2 Beyond worst-case analysis
In our updated histogram step, we add Laplace noise with scale $2/[n\varepsilon]$ to the entries of the (non-DP) NN histogram that are positive, and then threshold the noisy entries at $H=2\log(1/\delta)/(n\varepsilon)+1/n$. More concretely, let $\hat{\mu}=\operatorname{NN\_histogram}(S,V,\rho)$ be the non-DP histogram between a dataset $S$ and a set of variations $V\in\Omega^{m}$ (see Algorithm 4). The noisy histogram $\tilde{\mu}$ is then given by $\tilde{\mu}[i]=(\hat{\mu}[i]+L_{i})\mathbbm{1}_{(\hat{\mu}[i]>0,\hat{\mu}[i]+L_{i}\geq H)}$ where $\{L_{i}\}_{i\in[m]}\overset{iid}{\sim}\text{Lap}(2/n\varepsilon)$. While this algorithm was proven to be $(\varepsilon,\delta)$-DP in Theorem $3.5$ of [39], to the best of our knowledge, its utility guarantee in terms of $W_{1}$ distance is new. Its proof can be found in Appendix G.
@@
To demonstrate when this proposition yields a tighter result, consider the case when $|\tilde{I}|$ and $|\hat{I}|$ are small. As an extreme example, suppose $S$ consists of one data point repeated $n$ times. Then, $|\tilde{I}|=1$ and $|\hat{I}|=0$, leading to an error of $O(\frac{D\log(1/\delta)}{n\varepsilon}+\frac{D}{n})$. A similar intuition is valid for highly clustered datasets. On the other extreme, if every positive entry of $\hat{\mu}$ is $1/n$, then $|\tilde{I}|=|\hat{I}|=n$, leading to a trivial error bound. See Appendix E for simulations.
@@ 5 Connections between PE and Private Signed Measure Mechanism
In this section, we show connections between PE and the Private Signed Measure Mechanism (PSMM), a DP synthetic data method with formal $W_{1}$ guarantees from [24]. Despite its theoretical guarantees, PSMM is not currently used for high-dimensional, real datasets.
@@
PSMM takes a dataset $S$ and a partition $\{\Omega_{i}\}_{i\in[m]}$ of $\Omega$. It privately counts how many points fall into each bin $\Omega_{i}$, and then builds a DP distribution over $\Omega$ proportional to these counts (details in Appendix H). Synthetic data is sampled from this distribution.
We can reinterpret PSMM in the language of our framework. Instead of a partition, we can provide a discrete support $\Omega^{\prime}\subset\Omega^{m}$ such that the Voronoi partition induced by $\Omega^{\prime}$ coincides with $\{\Omega_{i}\}_{i\in[m]}$. Counting how many elements fall into each $\Omega_{i}$ is then equivalent to counting how many points in $S$ have $\Omega^{\prime}[i]$ as their nearest neighbor. The resulting private distribution can thus be viewed as a DP nearest-neighbor histogram over $\Omega^{\prime}$. We believe that the reason behind this (modified PSMM) algorithm working is that the NN histogram solves a 1-Wasserstein minimization problem (Proposition 5.1; proof in Appendix G), a property that also plays a key role in our analysis of PE.
@@
6 Simulations
We now show how our theory has the potential to explain the behavior of PE in practice. We do so with simulations, where we consider the sample space $\Omega$ to be the unit $\ell_{2}$ ball in $\mathbb{R}^{2}$ and the sensitive dataset is random over $\Omega\cap\mathbb{R}^{2}_{+}$. We use privacy parameters $\varepsilon=1,\delta=10^{-4}$. Our theory indicates that $O(\log(n\varepsilon))$ PE steps ensure convergence; for experiments we set $T=2\log(n\varepsilon)$. The noise $\sigma$ is then computed with the analytic Gaussian mechanism [4, Theorem 8]. The rest of the parameters are set as indicated in Theorem 4.1. We post-process noisy histograms by truncating at $H=0$ and then re-normalizing.
Figure 2: The best number of PE steps depends heavily on the initialization from $\operatorname{Random\_API}$. The private dataset $S$ is a random dataset in $\Omega\cap\mathbb{R}^{2}_{+}$ Left: When $S_{0}=S$ (i.e., $\Gamma_{0}=0$, the private data is the same as the initial synthetic data), the optimal number of PE steps is 0. Middle: When PE is initialized poorly (e.g., $S_{0}$ consists of only $(0,0)$, so $\Gamma_{0}$ is large), more iterations are needed. Right: Interpolatating between the previous cases parametrized by $\beta$: $S_{0}=(1-2\beta)S$, PE can improve or degrade performance depending on $\operatorname{Random\_API}$; it never exceeds the worst-case error bound. Results are averaged over 100 runs. Our analysis explains this phenomenon (see text).
Our analysis explains dependence on the initialization.
Let $\Gamma_{t}\triangleq\mathbb{E}[W_{1}(\mu_{S},\mu_{S_{t}})]$, i.e. the expected Wasserstein-1 distance between the private data and the $t$th iteration of synthetic data. From equation (2) in the proof of Theorem 4.1, we obtain $\Gamma_{T}\leq(1-\gamma)^{T}(\Gamma_{0}-err)+err$ where $err$ is an error term vanishing with the number of samples. We note from this equation that if $\Gamma_{0}\leq err$, then the optimal number of steps for PE is $T=0$. However, when $\Gamma_{0}$ is large, our worst-case upper bound will be more accurate. Figure 2 illustrates this phenomenon.
Our findings align with prior empirical work. [31] found that PE requires more steps to converge and consistently improves over time when $\Gamma_{0}$ is large. In addition, in [38], it was found that in many instances PE worsens with time, and the optimal number of steps is $1$. Our theory shows how both of these cases are indeed possible. This observation motivates the need for an analysis of PE beyond the worst case, which we briefly addressed in Section 4.2.
Our theory can guide the practice of PE.
Our theory suggests certain parameter setting for PE. For example, one takeaway from Theorem 4.1 is that setting the number of PE steps $T=O(\log(n\varepsilon/\sqrt{\log(1/\delta)}))$ (independent of $d$) suffices to ensure convergence under $(\varepsilon,\delta)$-DP. Another is that the number of synthetic samples must be chosen to balance the error due to adding noise to the NN histogram in Step 7 (which increases with number of synthetic samples) with the error due to sampling with replacement in Step 9 (which decreases with the number of synthetic samples) of Algorithm 2. Figure 3 shows how picking the number of PE steps or the number of synthetic data points suboptimally can hurt the final accuracy.

### v2

@@ 4.1 Convergence of PE in Euclidean space
We start by presenting the details of the algorithm that we analyze. Consider a metric (data) space $(\Omega,\rho)$ and a private dataset $S\in\Omega^{n}$. Even though we will use the notation $(\Omega,\rho)$ throughout this section, we assume in this subsection that $\Omega\subset\mathbb{R}^{d}$ is a convex and compact with $\operatorname{diam}(\Omega)\leq D$ and $\rho(\cdot,\cdot)$ is the $\ell_{2}$ distance. We also assume Algorithm 2 uses the same APIs as Algorithm 1, as described in Section 3.
Algorithm 2 Private Evolution with $D_{BL}$ projection; steps in blue differ from practical PE PE23
@@
Algorithm 2 iteratively refines an initial random dataset $S_{0}$ using DP nearest-neighbor information. At each iteration, it generates variations $V_{t}$ from the current synthetic dataset $S_{t-1}$ and computes a histogram $\hat{\mu}_{t}$ indicating how often each candidate is the nearest neighbor of points in the sensitive dataset $S$. Gaussian noise is added to this histogram to ensure differential privacy, resulting in a noisy vector $\tilde{\mu}_{t}$. Rather than using thresholding like Algorithm 1 in Line 8, this algorithm projects $\tilde{\mu}_{t}$ onto the probability simplex using the $D_{BL}$ distance in Line 8, yielding a valid distribution $\mu_{t}^{\prime}$ over $V_{t}$ that remains close to the noisy estimate under the BL metric. The next synthetic dataset $S_{t}$ is created randomly by sampling $n_{s}$ points from $V_{t}$ according to $\mu_{t}^{\prime}$ in Line 9, as opposed to Algorithm 1 which creates $S_{t}$ deterministically in Line 9. This process is repeated for $T$ iterations, after which the final synthetic dataset $S_{T}$ is returned. Note that unlike Algorithm 1, our version receives the number of synthetic data points as input. We provide a comparison between Algorithms 1 and 2 in Appendix C.
@@
Algorithm 2 is nearly identical to the practical PE implementation in PE23, differing only in how the noisy histogram $\tilde{\mu}_{t}$ is post-processed (highlighted in blue). We project $\tilde{\mu}_{t}$—a signed measure from Gaussian noise added to a histogram—onto $\mathcal{P}(V_{t})$ by minimizing the bounded Lipschitz distance, which can be done via linear programming (e.g., Algorithm 2 from Vershynin23DPsyntheticdata). This is crucial for obtaining worst-case convergence guarantees. In contrast, PE23 applies a simpler heuristic: thresholding $\tilde{\mu}_{t}$ at some $H>0$ and re-normalizing, which is more efficient but can discard all votes if values fall below the threshold, making it unsuitable for worst-case analysis. In Section 4.2, we provide data-dependent accuracy bounds for a related noisy histogram obtained from adding Laplace noise and thresholding. These bounds are vacuous in the worst case but can be much tighter in favorable regimes, e.g., if the private data is highly clustered.
@@
We are now ready to present our main result: an upper bound on expected utility of Algorithm 2. We provide a proof sketch, and the full proof is in Appendix E.
@@
Let $(\Omega,\|\cdot\|_{2})$ with $\Omega\subset\mathbb{R}^{d}$ closed and convex and $\operatorname{diam}(\Omega)\leq D$ be the sample space. Fix $\sigma\in(0,1)$ and let $\gamma=\frac{1}{8\pi[(\sqrt{d}+\log(2))^{2}+\log(2)]}$. Suppose $\alpha=D\sigma^{1/\max\{d,2\}}$ in (1), and let $S_{T}$ be the output of Algorithm 2 run on input $S\in\Omega^{n}$$T\geq\bigg\lceil\log\bigg(\frac{\gamma}{D\sigma^{1/\max\{d,2\}}}\bigg)\big/\gamma\bigg\rceil,n_{s}=\sigma^{-1}(2\lceil\log_{2}(D/\alpha)\rceil+1)^{1/\max\{d,2\}-1}$, $\sigma,\|\cdot\|_{2}$. Then
@@
Furthermore, if $\varepsilon,\delta\in(0,1)$, and we instead run Algorithm 2 with the same parameter setting as above except $T=\bigg\lceil\log\bigg(\frac{\gamma(n\varepsilon)^{1/\max\{d,2\}}}{\big(4\sqrt{\log(1/\delta)}\big)^{1/\max\{d,2\}}}\bigg)\big/\gamma\bigg\rceil$ and $\sigma=\frac{4\sqrt{T\log(1.25/\delta)}}{n\varepsilon}$, then the algorithm is $(\varepsilon,\delta)$-DP and its output $S_{T}$ satisfies
@@
Some comments are in order. First, note that $n_{s}\geq 1$ for $\sigma\in(0,1)$ and to ensure $\sigma\in(0,1)$ under DP we only need $\sqrt{T\log(1/\delta)}\lesssim n\varepsilon$ . Second, our analysis suggests that we should set number of evolution steps $T=\tilde{O}(\log(n\varepsilon))$ and the number of synthetic samples $n_{s}=\tilde{O}(n\varepsilon/\sqrt{T})$, which we use in Section 6.333In the nonprivate setting $(\epsilon=\infty)$, one can skip the noise addition and sampling steps and set $n_{s}=n$, since $n_{s}$ does not need to be chosen to balance the error terms that arise from DP. This results in a nonprivate algorithm similar to that of PE23 with an expected $W_{1}$ error of $2\alpha$ after $T\gtrsim d\log(D/\alpha)$ evolution steps without DP constraints, similar to Theorem $1$ of PE23. Finally, our rates are vacuous for $d\gtrsim\log(n)$. This is expected when using $W_{1}$ as a utility metric, as evidenced by the lower bound of $n^{-1/d}$ under pure DP from (boedihardjo2024private, Theorem 8.5).
@@
At a high level, the convergence of PE follows from two facts:
@@
These statements are made rigorous in our proof. Regarding the first statement, Lemmata E.1 and E.2 give, $\mathbb{E}_{t}[W_{1}(\mu_{S},\hat{\mu}_{t+1})]\leq(1-\gamma)W_{1}(\mu_{S},\mu_{S_{t}})+\alpha$ for some $\gamma=\Theta(1/d)$, where $\mathbb{E}_{t}$ denotes the expectation conditioned on the randomness of the algorithm up to iteration $t$.
To bound $D_{BL}(\hat{\mu}_{t+1},\tilde{\mu}_{t+1})$, we use the fact that $\tilde{\mu}_{t+1}=\hat{\mu}_{t+1}+Z$ with $Z\sim\mathcal{N}(0,\sigma^{2}I_{|V_{t}|})$, which by definition implies $D_{BL}(\hat{\mu}_{t+1},\tilde{\mu}_{t+1})=\sup_{f\in{\cal F}_{BL}}\int_{\Omega}f(d\hat{\mu}_{t+1}-d(\hat{\mu}_{t+1}+Z))=\sup_{f\in{\cal F}_{BL}}\sum_{i\in[|V_{t+1}|]}f(V_{t+1}[i])Z_{i}$, which is the supremum of a Gaussian process. The expectation of this term can be bounded using empirical process theory; Lemma E.3 states that $\mathbb{E}_{t}[D_{BL}(\hat{\mu}_{t+1},\tilde{\mu}_{t+1})]\leq|V_{t}|\sigma G_{|V_{t}|}({\cal F}_{BL})$, where $G_{|V_{t}|}({\cal F}_{BL})$ is a Gaussian complexity term bartlett2002rademacher. This term can be bounded using Corollary E.1, which is a consequence of Lemma E.4. We use Dudley’s chaining to prove these results. We remark that our technique prove this bound resembles Vershynin23DPsyntheticdata, with the difference that they deal with a Laplacian complexity arising from the analysis of a pure-DP algorithm that uses the Laplace mechanism dwork2014DPfoundations.
Finally, to control $W_{1}(\mu^{\prime}_{t+1},\mu_{S_{t+1}})$ we use results from the literature of the convergence of empirical measures in the Wasserstein distance lei2020empiricalwasserstein; fournier2015empiricalwasserstein that quantify the Wasserstein distance between a measure and the empirical measure of iid samples from it; see Lemma E.5. We obtain $\mathbb{E}_{t}[W_{1}(\mu^{\prime}_{t+1},\mu_{S_{t+1}})]\leq\tilde{O}\big(Dn_{s}^{-1/\max\{2,d\}}\big)$. Putting all the inequalities together, we conclude that
$\mathbb{E}_{t}[W_{1}(\mu_{S},\mu_{S_{t+1}})]\leq(1-\gamma)W_{1}(\mu_{S},\mu_{S_{t}})+\alpha+2|V_{t}|\sigma G_{|V_{t}|}({\cal F}_{BL})+\tilde{O}\big(Dn_{s}^{-1/\max\{2,d\}}\big).$
Recursing this inequality, we can get a bound on $\mathbb{E}[W_{1}(\mu_{S},\mu_{S_{T}})]$ while carefully balancing the number of variations with the number of synthetic samples so that the error terms are as small as possible. This finishes the proof of $\mathbb{E}\left[W_{1}(\mu_{S},\mu_{S_{T}})\right]\leq\tilde{O}(dD\sigma^{1/d})$.
Regarding the second result, for privacy we use the fact that $\sigma=\frac{\sqrt{4T\log(1.25/\delta)}}{n\varepsilon}$ is enough to ensure $(\varepsilon,\delta)$-DP of an adaptive composition of $T$ Gaussian mechanisms where each of them adds noise to a NN histogram with $\ell_{2}$ sensitivity $\sqrt{2}/n$ (dong2022gaussianDP, Corollary 3.3). For convergence, we plug in the expression for $T$ in $\sigma=\frac{\sqrt{4T\log(1.25/\delta)}}{n\varepsilon}$ and then $\sigma$ in $\mathbb{E}\left[W_{1}(\mu_{S},\mu_{S_{T}})\right]\leq\tilde{O}(d\sigma^{1/d})$. ∎
@@
Most of our proof techniques work in metric spaces. Under some conditions on the APIs, the analysis of Algorithm 2 can be extended to more general Banach spaces (Appendix G).
@@ 4.2 Beyond worst-case analysis
In our updated histogram step, we add Laplace noise with scale $2/[n\varepsilon]$ to the entries of the (non-DP) NN histogram that are positive, and then threshold the noisy entries at $H=2\log(1/\delta)/(n\varepsilon)+1/n$. More concretely, let $\hat{\mu}=\operatorname{NN\_histogram}(S,V,\rho)$ be the non-DP histogram between a dataset $S$ and a set of variations $V\in\Omega^{m}$ (see Algorithm 4). The noisy histogram $\tilde{\mu}$ is then given by $\tilde{\mu}[i]=(\hat{\mu}[i]+L_{i})\mathbbm{1}_{(\hat{\mu}[i]>0,\hat{\mu}[i]+L_{i}\geq H)}$ where $\{L_{i}\}_{i\in[m]}\overset{iid}{\sim}\text{Lap}(2/n\varepsilon)$. While this algorithm was proven to be $(\varepsilon,\delta)$-DP in Theorem $3.5$ of Vadhan2017complexityofDP, to the best of our knowledge, its utility guarantee in terms of $W_{1}$ distance is new. Its proof can be found in Appendix H.
@@
[…]
@@ 5 Connections between PE and Private Signed Measure Mechanism
In this section, we show connections between PE and the Private Signed Measure Mechanism (PSMM), a DP synthetic data method with formal $W_{1}$ guarantees from Vershynin23DPsyntheticdata. Despite its theoretical guarantees, PSMM is not currently used for high-dimensional, real datasets.
@@
PSMM takes a dataset $S$ and a partition $\{\Omega_{i}\}_{i\in[m]}$ of $\Omega$. It privately counts how many points fall into each bin $\Omega_{i}$, and then builds a DP distribution over $\Omega$ proportional to these counts (details in Appendix I). Synthetic data is sampled from this distribution.
We can reinterpret PSMM in the language of our framework. Instead of a partition, we can provide a discrete support $\Omega^{\prime}\subset\Omega^{m}$ such that the Voronoi partition induced by $\Omega^{\prime}$ coincides with $\{\Omega_{i}\}_{i\in[m]}$. Counting how many elements fall into each $\Omega_{i}$ is then equivalent to counting how many points in $S$ have $\Omega^{\prime}[i]$ as their nearest neighbor. The resulting private distribution can thus be viewed as a DP nearest-neighbor histogram over $\Omega^{\prime}$. We believe that the reason behind this (modified PSMM) algorithm working is that the NN histogram solves a 1-Wasserstein minimization problem (Proposition 5.1; proof in Appendix H), a property that also plays a key role in our analysis of PE.
@@
6 Empirical Results
Simulations
We begin with simulations where the sample space $\Omega$ is the unit $\ell_{2}$ ball in $\mathbb{R}^{2}$ and the sensitive dataset is random over $\Omega\cap\mathbb{R}^{2}_{+}$. We use privacy parameters $\varepsilon=1,\delta=10^{-4}$. Our theory indicates that $O(\log(n\varepsilon))$ PE steps ensure convergence; for simulations we set $T=2\log(n\varepsilon)$. The noise $\sigma$ is then computed with the analytic Gaussian mechanism (balle2018improvingGaussianmech, Theorem 8). The remaining parameters are set as in Theorem 4.1. We post-process noisy histograms by truncating at $H=0$ and then re-normalizing.
Dependence on initialization.
Let $\Gamma_{t}\triangleq\mathbb{E}[W_{1}(\mu_{S},\mu_{S_{t}})]$, i.e. the expected Wasserstein-1 distance between the private data and the $t$th iteration of synthetic data. From equation (2) in the proof of Theorem 4.1, we obtain $\Gamma_{T}\leq(1-\gamma)^{T}(\Gamma_{0}-err)+err$ where errerr increases polynomially with $T$. We note from this equation that if $\Gamma_{0}\leq err$, then the optimal number of steps for PE is $T=0$. However, when $\Gamma_{0}$ is large, our worst-case upper bound will be more accurate. Figure 5 illustrates this phenomenon. Our findings help to explain prior empirical work. PE23 found that PE requires more steps to converge and consistently improves over time when $\Gamma_{0}$ is large. In contrast, swanberg2025PEtabulardata found that in many instances, PE worsens with time, and the optimal number of steps is $1$.
Hyperparameter selection.
Theorem 4.1 suggests setting the number of PE steps $T=O(\log(n\varepsilon/\sqrt{\log(1/\delta)}))$ (independent of $d$). Similarly, the number of synthetic samples $n_{s}$ should be chosen to balance the error due to adding noise to the NN histogram in Step 7 of Algorithm 2 (which increases with number of synthetic samples) with the error due to sampling with replacement in Step 9 (which decreases with the number of synthetic samples). Figure 2 shows how a suboptimal number of PE steps or synthetic data points degrades data quality.